Shift Summary - Evening - 28 June 1999

Douglas A. Jensen (djensen@fnal.gov)
Tue, 29 Jun 1999 00:06:28 -0500 (CDT)

Shift Summary for evening shift, 28/June/99

On Shift: Jim Graham, Doug Jensen

And, Hogan, Rick K, Sasha, Ed (a while) ...

Hours of beam delivered at all intensities: 8
Hours of beam delivered at nominal intensity: 1.5
Hours of E832 data (special runs not included): 0.7
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:changing tapes, accelerator down
Accelerator problems, restart KTeV2, assorted front end problems.....

========================================================================
Runs ended this shift:
Run 13753 Beam_832 much beam tuning, ended with KTeV2 crashing
Run 13754 Beam_832 Aborted for drive problem
Run 13755 Beam_832 aborted....

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Took over running experiment, but accelerator tuning. Running 5 batches
Beam quality ???l Boa 1.47
6pm still tweeking beam. Spike at beginning. Intensity slowily
increasing.
10:10 MCR finds problems with TeV aborts. Returnes to get to about
8E12 with 11 batches. Beam nice.
But an assortment of problems:
Tape drive stuck in drive
KTeV2 crashes, requires reboot
Cables supplying ordinal number reseated....
Several dart restarts ... !

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13753 15:30 Running, new shift crew taking over
Note that no gas check list done, no new He trailer
160 16:00 Getting beam. debate quality, still tuning, 5 bunches
250 18:00 Beam tuning continues. See spills to 5.2E12.
Big spikes at start of beam. BOA about 1.4
DN swaps tubes on He trailer.
Alcohol temp about 4 deg C.
305 19:00 MCR calls, they are going back to 5e12 and 5 batches
to try to understand abort problems.
403 21:20 Run ended with system crash.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift